OpenStudy (mkmkasim):

antiderivative of ∫(1+x)/(1+x^2).

OpenStudy (anonymous):

try \(u=1+x^2\)

OpenStudy (anonymous):

no, that is a bad idea

Parth (parthkohli):

^ Why?

OpenStudy (anonymous):

try \[\frac{1}{1+x^2}+\frac{x}{1+x^2}\] that will work nicely

OpenStudy (anonymous):

first is one you do with your eyeballs, (i.e. one you have memorized) second is a mental u - sub

Parth (parthkohli):

Oh, I see

OpenStudy (turingtest):

sorry, @mkmkasim that was an accident :S

OpenStudy (mkmkasim):

is it arctanx + ln|1+x^2| +c

OpenStudy (anonymous):

first is right, second you are off by a constant

OpenStudy (mkmkasim):

how?

OpenStudy (anonymous):

\(u=1+x^2,\frac{1}{2}du=xdx\)

OpenStudy (anonymous):

need a \(\frac{1}{2}\) out front to adjust for the u sub

OpenStudy (mkmkasim):

oh okay I got it. Thanks